### v1

# Scalable Unit Harmonization in Medical Informatics Using Bi-directional Transformers and Bayesian-Optimized BM25 and Sentence Embedding Retrieval

Abstract
Objective: To develop and evaluate a scalable methodology for harmonizing inconsistent units in large-scale clinical datasets, addressing a key barrier to data interoperability.
Materials and Methods: We designed a novel unit harmonization system combining BM25, sentence embeddings, Bayesian optimization, and a bidirectional transformer based binary classifier for retrieving and matching laboratory test entries. The system was evaluated using the Optum Clinformatics Datamart dataset (7.5 billion entries). We implemented a multi-stage pipeline: filtering, identification, harmonization proposal generation, automated re-ranking, and manual validation. Performance was assessed using Mean Reciprocal Rank (MRR) and other standard information retrieval metrics.
Results: Our hybrid retrieval approach combining BM25 and sentence embeddings (MRR: 0.8833) significantly outperformed both lexical-only (MRR: 0.7985) and embedding-only (MRR: 0.5277) approaches. The transformer-based reranker further improved performance (absolute MRR improvement: 0.10), bringing the final system MRR to 0.9833. The system achieved 83.39% precision at rank 1 and 94.66% recall at rank 5.
Discussion: The hybrid architecture effectively leverages the complementary strengths of lexical and semantic approaches. The reranker addresses cases where initial retrieval components make errors due to complex semantic relationships in medical terminology.
Conclusion: Our framework provides an efficient, scalable solution for unit harmonization in clinical datasets, reducing manual effort while improving accuracy. Once harmonized, data can be reused seamlessly in different analyses, ensuring consistency across healthcare systems and enabling more reliable multi-institutional studies and meta-analyses.
1 Background and Significance
Modern scientific research increasingly depends on the integration of heterogeneous data sources, with inconsistent units of measurement posing a persistent challenge, particularly in laboratory analyses where identical quantities may be reported using diverse conventions. Data harmonization is a critical prerequisite for large-scale clinical research, ensuring interoperability across electronic health record systems and supporting reliable, reproducible medical analyses. Laboratory unit harmonization is especially complex due to domain-specific terminology and the clinical importance of precise unit conversion. Traditional methods for unit harmonization often rely on manually curated rule-based systems and mapping tables developed by domain experts [1]. While these approaches can be effective within narrowly defined contexts, they are not easily scalable and require ongoing maintenance to accommodate emerging laboratory codes and units. This labor-intensive process introduces bottlenecks in research workflows, delays analysis, and increases resource burdens. Emerging methodologies that combine machine learning with information retrieval techniques offer promising solutions to these challenges. Techniques such as sentence embeddings, bidirectional transformers, and optimized retrieval models have demonstrated potential for automating and scaling harmonization processes with improved accuracy and efficiency [2, 3, 4]. In this paper, we present a novel automated unit harmonization system that advances current practices by integrating Bayesian-optimized BM25, sentence embeddings, and transformer-based re-ranking mechanisms. Our system directly addresses major limitations in existing frameworks, including poor scalability to billions of records, limited adaptability to diverse naming conventions, inadequate contextual understanding of laboratory codes, lack of dynamic feedback loops, and overreliance on manual validation. By mitigating these constraints, our approach facilitates seamless data integration, reduces error rates, and accelerates the pace of scientific discovery, ultimately empowering clinical researchers to derive more robust, timely, and reproducible insights from increasingly complex and voluminous healthcare datasets.
2 Materials and Methods
2.1 Datasets
2.1.1 Primary Dataset
The Optum Clinformatics Data Mart (CDM) constitutes our primary evaluation dataset. This comprehensive, de-identified database integrates administrative health claims data from over 84 million individuals across all 50 U.S. states. The resource contains more than 7.5 billion medical and pharmacy claims, documenting healthcare utilization and associated costs. Data elements include member demographics, detailed medical and pharmacy claims, laboratory results, inpatient confinement records, and provider information.
In this study, we used the 2024 Q1, Q2, and Q3 version of the Optum Clinformatics Data Mart. To support the harmonization process, key fields were extracted from the dataset, including LOINC codes, units of measurement, frequency of occurrence, and descriptive statistics such as minimum, maximum, mean, and standard deviation. This process resulted in approximately 30,000 unique triads (test, sample, unit) requiring standardization. Of these, 17,244 entries were identified in our internal database as candidates for matching.
2.1.2 Reference Dataset
We utilized an internal Labcodes Standard Database for unit harmonization, providing a comprehensive mapping of laboratory tests with standardized information. The database facilitates consistent data representation across diverse sources by linking original units to preferred units and providing necessary conversion factors.
The harmonization process is mapped against multiple fields within this database, including Test Name, Test Label, Synonym, Sample Name, Labcode, Preferred Unit, Actual Unit, Conversion Factor, and various CDISC standardization fields. Although all fields contribute to a comprehensive test definition, harmonization can often be achieved with a minimum of Test Name, Sample Name, and Unit. Inclusion of additional fields improves the accuracy and specificity of the match.
2.2 System Architecture
Our harmonization framework implements a multi-stage pipeline comprising data processing, predictive modeling, and validation components. A key design principle is the modular decomposition of this framework into independent blocks, allowing for focused development, rigorous testing, and iterative improvement of individual components.
2.2.1 Overall Pipeline Structure
The overall system is structured as an end-to-end processing pipeline (Fig. 1) with distinct stages, each designed to address specific challenges in the harmonization process. The pipeline begins with the Download stage, which involves the extraction of data from the relational database management system (RDBMS), focusing on unique combinations of LOINC codes and reported units, accompanied by pre-computed usage statistics. Following this, the Filtering stage removes entries that fall below a predefined frequency threshold. During this phase, the system also enforces unit name normalization and eliminates inconsistencies to ensure data quality and reduce noise in downstream tasks.
The Update stage enriches the filtered dataset by incorporating additional contextual information. This is achieved through inner joins using LOINC codes to retrieve descriptive labels and standardized mappings for column names, enhancing interpretability and facilitating accurate retrieval. Next, the Predictive Modeling stage implements the core harmonization intelligence (detailed in Section 2.2.2). The pipeline concludes with the Manual Validation stage, which allows domain experts to review and confirm harmonization results through a custom-built user interface. This feedback not only ensures clinical safety and accuracy but also contributes to the iterative improvement of the system.
Figure 1: Overview of the harmonization pipeline showing the sequential flow of data from raw database entries to validated harmonized units
2.2.2 Predictive Model Architecture
The predictive model architecture (Fig. 2) constitutes the core intelligence of our system and consists of three primary components. The first component is the Hybrid Retrieval Engine, which combines two complementary approaches. The Lexical Retrieval Module implements BM25 algorithm-based traditional term-matching for precise keyword identification and excels at exact and fuzzy matching of laboratory terminology. Complementing this, the Semantic Retrieval Module utilizes sentence embeddings to capture contextual relationships between terminology and can identify semantically equivalent terms even when they share limited lexical overlap.
The second component is the Bayesian Score Optimizer, which adjusts the weights between lexical and semantic scores to maximize retrieval performance. This component uses a Gaussian Process-based optimization procedure to determine the optimal balance between precision and recall for each query context. Importantly, this optimization is performed offline, prior to inference, ensuring that the optimal weights are precomputed and ready for use during query execution.
The third component is the Transformer-based Reranker, a bidirectional transformer model that acts as the final decision layer, evaluating and reordering candidate matches based on deeper contextual understanding. This component addresses cases where the initial retrieval may have missed subtle semantic distinctions, particularly in complex medical terminology where minor variations can signify clinically significant differences.
Figure 2: Components of the predictive model architecture showing the interaction between retrieval mechanisms and contextual reranking
The retrieval engine is implemented using an Elasticsearch backend, supporting flexible and scalable querying through JSON-based schema definitions which govern the indexing and searching of laboratory unit metadata. The hybrid retrieval strategy combines lexical scoring—based on BM25 metrics—with semantic similarity, computed via cosine distance between sentence embeddings. The final ranking score for each candidate is calculated as a weighted linear combination of these two components, with weights determined by the Bayesian optimizer.
This modular architecture enables both high accuracy and scalability. The retrieval components provide broad coverage of potential matches, while the reranker delivers precision by applying deeper semantic understanding. Each component can be independently optimized and updated as techniques advance, ensuring the system remains state-of-the-art while maintaining backward compatibility with existing data pipelines.
2.2.3 Sentence Embeddings
We employ the pre-trained transformer-based model bijaygurung/stella-en-400Mv5 [5] with a vector size of 1024 to generate dense vector representations for laboratory test descriptions. This model has been trained using Matryoshka Representation Learning (MRL) [6] to produce hierarchical embeddings at multiple granularities.
We also explored fine-tuning the sentence embedding on laboratory-specific terminology using a synthetic dataset of pairs labeled as either equivalent or dissimilar. However, our experiments demonstrated that the general-purpose model significantly outperformed the fine-tuned models.
2.2.4 BM25 Enhancement
We enhanced our retrieval system by incorporating the BM25 algorithm [7], a state-of-the-art probabilistic relevance framework. BM25 offers a more refined modeling of term frequency, along with document length normalization to mitigate frequency bias. The BM25 score for a document $d$ with respect to a query $q$ is computed as:
$\text{score}(d,q)=\sum_{t\in q}\text{IDF}(t)\cdot\frac{\text{TF}(t,d)\cdot(k_{1}+1)}{\text{TF}(t,d)+k_{1}\cdot\left(1-b+b\cdot\frac{|d|}{\text{avgdl}}\right)}$
where:
  • $\text{TF}(t,d)$ is the term frequency of term $t$ in document $d$
  • $\text{IDF}(t)$ is the inverse document frequency of term $t$
  • $|d|$ is the length of document $d$ in words
  • avgdl is the average document length in the collection
  • $k_{1}$ (typically 1.2-2.0) and $b$ (typically 0.75) are free parameters that control term frequency scaling and document length normalization respectively
Our system builds on the strengths of the BM25 ranking function to support both exact and fuzzy matching, optimizing retrieval performance across varying levels of lexical precision. In scenarios requiring exact matching, BM25 prioritizes documents that contain query terms with higher frequency and in close proximity. This capability is particularly beneficial in the context of laboratory code harmonization, where precise terminology—such as LOINC codes or standardized unit expressions—is critical for accurate identification and mapping.
In addition to exact matches, the system supports fuzzy matching to accommodate natural variations in terminology, including differences in spelling, formatting, or word choice. BM25 introduces a graded penalty for such partial matches, allowing relevant documents with similar but non-identical terms to be retrieved while preserving ranking quality. This is especially important in healthcare data, where naming conventions can vary across institutions or over time.
To further enhance retrieval effectiveness, the system incorporates curated synonym lists that expand the query space by mapping equivalent or related terms. This extension is particularly valuable in laboratory contexts, where a single concept may be referred to using multiple terminologies. By integrating synonym expansion into the BM25 framework, the system ensures comprehensive coverage and improves recall without sacrificing precision. Collectively, these enhancements enable the retrieval engine to support robust, flexible, and clinically meaningful harmonization of laboratory units.
2.2.5 Search and Ranking
Our hybrid search system integrates both lexical and semantic search capabilities through a multi-tiered (Fig. 2) designed to enhance retrieval accuracy and contextual relevance. The first component of this architecture involves the generation of combined text embeddings that represent laboratory test descriptions alongside relevant metadata. These embeddings enable the system to capture nuanced semantic relationships between queries and candidate entries, going beyond surface-level term matching.
To refine retrieval precision, the system employs field-specific search boosts that reflect the relative importance of different attributes. For instance, fields such as standardized test names or unit designations may carry more weight in the matching process than auxiliary metadata. These boost factors are configurable and allow the system to dynamically adjust the influence of each field depending on the context of the search.
Structured queries are constructed to leverage both exact and fuzzy matching mechanisms, alongside semantic vector similarity. The system utilizes a custom-built module that parses incoming queries using field-aware syntax, applies the predefined boost factors, and computes a composite score for each candidate result. This score is derived through a weighted combination of BM25 lexical scores and cosine similarity from semantic embeddings, allowing for a balanced integration of symbolic and contextual retrieval. By merging these complementary strategies, the system achieves high recall and precision in laboratory unit harmonization tasks, even in the presence of terminological variation and incomplete data.
2.2.6 Bayesian Optimization
To determine the optimal combination of retrieval methods, we implemented Bayesian optimization to fine-tune the weights assigned to various components of the search process. This approach allows for the systematic exploration of the search parameter space, ensuring that the retrieval system operates at peak efficiency while balancing competing factors. The optimization process primarily focused on adjusting field boost weights, which control the relative importance of different fields (e.g., test descriptions or units) in the retrieval process. By tuning these parameters, the system can adapt to the specific needs of each search context.
Another key aspect of the optimization was balancing the contributions from the lexical (BM25) and semantic (embedding-based) search components. Both approaches offer unique advantages: lexical search excels at exact term matching, while semantic search captures deeper contextual relationships. The optimization process sought to find the best weighting between these two components to maximize retrieval effectiveness across diverse queries and data types.
Finally, the optimization aimed to maximize the Mean Reciprocal Rank (MRR), a widely used performance metric that evaluates the ranking quality of retrieved results. By focusing on MRR, the system ensures that relevant harmonization candidates appear near the top of the result list, enhancing the user experience and enabling faster, more accurate decision-making in laboratory unit harmonization tasks. Through this optimization process, we were able to fine-tune the retrieval system to deliver the most relevant results while minimizing errors and inconsistencies.
2.3 Reranker
The reranker component of the system leverages a bidirectional transformer architecture based on RoBERTa [8] to classify and rank candidate harmonizations. Specifically, we employed the RoBERTa-base-PM-M3-Voc-train-longer model, which was pre-trained on PubMed and PMC articles as well as MIMIC-III clinical notes, using a byte-pair encoding vocabulary learned from PubMed with an additional 50,000 training steps. This domain-adapted model provides enhanced performance on biomedical terminology compared to general-purpose language models. We fine-tuned this biomedical RoBERTa variant on a synthetic dataset consisting of 3,135,557 balanced pairs of laboratory test triads (test, sample, unit), with each pair labeled as either equivalent (1) or non-equivalent (0). The classifier was trained to predict the probability of equivalence between input triads, with this probability score subsequently used as the differentiative ranking metric. This architecture enables the reranker to evaluate each candidate by comparing it directly with the input query, ensuring that the most relevant harmonization suggestions are ranked at the top.
The model also includes token handling mechanisms tailored to the unique terminology of laboratory tests. This is critical, as laboratory names and unit expressions often involve highly specialized vocabulary that can vary across different systems and institutions. By integrating these mechanisms, the reranker can more accurately process domain-specific terms and handle variations in naming conventions.
To further improve the model’s ability to differentiate between similar tests, the reranker incorporates contrastive learning objectives. This helps the system refine its distinction between closely related test descriptions, enabling it to make more nuanced decisions when harmonizing terms that appear almost identical but have subtle differences in meaning or context.
The model was trained on a synthetic dataset, which was created by pairing various test descriptions with different levels of dissimilarity between matching and non-matching pairs. This dataset simulates the real-world challenges of harmonizing laboratory data, with both easy and difficult matching cases. To fine-tune the model, cross-entropy loss was applied, allowing the system to learn to classify test pairs accurately across multiple laboratory categories.
An important feature of the system is its ability to allow for incremental improvement over time. As the system is deployed and interacts with real-world data, manually corrected values from operational use are incorporated into the training dataset. This process of continuous learning enables the model to progressively refine its understanding and handling of complex harmonization patterns, ensuring that it becomes more accurate and effective as it is used in practice.
2.4 Performance Evaluation
We evaluated the performance of our approach through a comprehensive suite of metrics, ensuring that both the retrieval and re-ranking components of the system were thoroughly assessed. For retrieval performance, we measured metrics such as Precision@k, Recall@k, and Mean Reciprocal Rank (MRR). These metrics allow us to evaluate how well the system retrieves relevant candidates across different ranks, ensuring that the most relevant harmonization suggestions appear early in the result set and assessing the system’s overall retrieval effectiveness.
The re-ranking performance was evaluated using standard classification metrics, including Accuracy, Precision, Recall, and the F1 score. These metrics provide insights into the effectiveness of the reranker in classifying and ranking candidate harmonizations. By analyzing these performance indicators, we could assess how well the reranker distinguishes between relevant and irrelevant harmonization proposals, and how balanced its classification performance is across different types of test descriptions.
To assess the scalability of the system, we measured key factors such as indexing throughput, query latency, batch processing efficiency, and memory usage. These metrics are critical for ensuring that the system can handle large datasets and operate efficiently at scale, especially when processing millions of laboratory records.
Additionally, the manual validation process was tracked through a custom tagging system, which categorized each harmonization decision into one of the following statuses: ”Missing,” ”Verified,” ”Pending,” ”Human,” ”Copy,” or ”Reranked”. This system enabled us to monitor the progress and quality of harmonization proposals, track user interactions with the system, and ensure that manual corrections were incorporated into the system’s learning process, contributing to its ongoing refinement.
3 Results
We evaluated our harmonization approach using a labeled dataset comprising 17,243 queries from the Optum database. All model parameters were optimized using Bayesian optimization on a stratified random subsample of 2,500 queries.
3.1 Comparative Approach Analysis
We conducted three controlled experiments to evaluate the performance of different retrieval paradigms, each designed to assess the strengths of various retrieval techniques.
The first experiment, Lexical-Only Retrieval, implemented a fully optimized lexical retrieval pipeline that combined BM25, field-specific weighting, fuzzy string matching, and synonym expansion. This approach focused solely on exact term matching and lexical features. The results from this experiment showed a best MRR of 0.7985, highlighting the effectiveness of lexical retrieval in terms of precision and matching the exact terms present in the query.
The second experiment, Embedding-Only Retrieval, evaluated the retrieval performance using exclusively sentence-level semantic similarity, derived from a general-purpose sentence encoder. This approach relied entirely on semantic embeddings to capture deeper contextual relationships between queries and candidate results. The best MRR in this case was 0.5277, demonstrating the ability of semantic embeddings to capture broader relationships but also revealing their limitations in terms of precise domain-specific matching.
The third experiment, Hybrid Retrieval, combined both semantic and lexical signals within a unified retrieval framework. This approach synthesized the strengths of both methods, using lexical signals for precision and embeddings for semantic coverage. The best MRR achieved by this hybrid approach was 0.8833, showcasing a significant improvement over the individual approaches. These results underscore the complementary nature of semantic and lexical retrieval, where embeddings enhance coverage and recall by capturing semantic relationships beyond exact keyword matching, while lexical signals provide precision and discriminative power for domain-specific terminology.
Together, these experiments demonstrate that a hybrid retrieval approach, which combines both semantic and lexical features, leads to the most effective retrieval performance, offering a balanced solution that maximizes both recall and precision.
3.2 Overall Performance Metrics
Table 1 presents comprehensive evaluation metrics for our hybrid retrieval model:
Table 1: Comprehensive evaluation metrics for the hybrid retrieval model
Metric Hybrid Model Value
Mean Reciprocal Rank (MRR) 0.8833
Mean Average Precision (MAP) 0.8131
Precision@10 0.5863
Recall@10 0.9700
NDCG@10 0.5291
Success@10 0.9700
MRR@10 0.8833
Queries Evaluated 17,243
Queries With Results 17,243 (100%)
The identical values of Recall@10 and Success@10 (both 0.9700) stem from our evaluation setup, which assumes exactly one relevant result per query.
3.3 Performance at Different Cutoff Thresholds
Table 2 shows retrieval metrics at varying rank cutoffs:
Table 2: Retrieval metrics at varying rank cutoffs
Metric k=1 k=3 k=5
Precision@k 0.8339 0.7116 0.6520
Recall@k 0.8339 0.9224 0.9466
NDCG@k 0.8339 0.5992 0.5353
Success@k 0.8339 0.9224 0.9466
MRR@k 0.8339 0.8746 0.8801
These metrics highlight the behavior of the retrieval system as we increase the cutoff threshold $k$:
- Recall@k and Success@k increase with larger $k$ Precision@k and NDCG@k decrease with increasing $k$ MRR@k increases slightly with larger $k$, eventually saturating
3.4 Reranker Performance
The transformer-based reranker was implemented as a post-processing stage after initial hybrid retrieval. The reranker selectively overrides the initial ranking when its confidence score exceeds that of the top-1 entry from the hybrid retrieval phase.
Absolute MRR Improvement: 0.10
This improvement represents a substantial enhancement over the already strong hybrid retrieval model, bringing the final system MRR to 0.9833. The reranker was particularly effective at correcting cases where lexically similar but semantically different terms were initially ranked higher than the correct match.
4 Discussion
The experimental results provide strong evidence for the effectiveness of our hybrid harmonization approach in medical terminology retrieval. The substantial performance gap between the hybrid model (MRR: 0.8833) and both the lexical-only (MRR: 0.7985) and embedding-only (MRR: 0.5277) variants confirms that these retrieval paradigms capture complementary aspects of query-document relevance.
The relatively poor performance of the embedding-only approach suggests that while general-purpose sentence encoders capture semantic relatedness, they may lack the specificity required for precise medical terminology matching. Conversely, the lexical approach demonstrates strong discriminative power but may miss semantically equivalent expressions with limited lexical overlap.
Our hybrid architecture effectively addresses these limitations by leveraging the complementary strengths of both approaches. The performance metrics across different rank cutoffs demonstrate that the system achieves an optimal balance between precision and recall, with over 83% of queries returning the correct result in the top position.
The addition of the transformer-based reranker as a final stage provides a critical refinement layer, addressing cases where the initial retrieval components make errors due to complex semantic relationships or domain-specific nuances in medical terminology. The significant improvement in MRR achieved by the reranker underscores the value of deep contextual understanding in medical term harmonization tasks.
Existing harmonization systems encounter several critical challenges that hinder their effectiveness in real-world clinical settings. One major limitation is the limited scalability of these systems, making it difficult to process the vast volumes of laboratory records required for large-scale clinical research. As the volume of data continues to grow, traditional methods struggle to handle the increasing demand for faster and more efficient harmonization.
Another significant challenge is the insufficient handling of variations in naming conventions and abbreviations. Laboratory codes and terminologies often vary across different institutions and datasets, and existing systems are not always equipped to manage these discrepancies effectively. This leads to inconsistencies in data representation, making it harder to perform accurate and reliable harmonization.
Furthermore, many harmonization systems fail to incorporate contextual information when matching laboratory codes. Without context, these systems may misinterpret or incorrectly align terms that appear similar but have different meanings depending on the context, which compromises the quality and accuracy of the harmonization process.
A related issue is the lack of efficient feedback mechanisms for continuous performance improvement. As laboratory terminologies evolve and new codes emerge, existing systems often do not have an effective way to incorporate feedback or adapt to changes in real-time, leading to stagnation and a failure to keep pace with evolving practices.
Finally, there is a high dependency on manual curation and validation in many current harmonization approaches. Manual intervention is time-consuming, prone to human error, and often not scalable, making it difficult to maintain quality and consistency as datasets grow. These challenges highlight the need for more automated, scalable, and adaptive harmonization systems in clinical settings.
Our system addresses these limitations through its modular architecture, hybrid retrieval approach, and reranker component. The Bayesian optimization of weightings between lexical and semantic components allows the system to adapt to the specific characteristics of medical terminology, while the reranker provides deeper contextual understanding.
The system’s performance on the large-scale Optum dataset (7.5 billion entries) demonstrates its scalability and effectiveness in real-world settings. By automating much of the harmonization process, it reduces the manual effort required while maintaining high accuracy.
5 Conclusion
We presented a scalable and efficient framework for unit harmonization in clinical datasets using a combination of BM25, sentence embeddings, a reranker based on bidirectional transformers, and Bayesian optimization techniques. Our system automates the harmonization process, reducing manual effort and improving accuracy. The results obtained from the Optum Clinformatics Data Mart dataset demonstrate that the methodology is effective and adaptable to large datasets, making it a promising solution for future healthcare data standardization efforts.
The implications of this work extend beyond technical achievements to address fundamental challenges in healthcare data management. By providing a consistent and standardized approach to unit harmonization, our framework significantly enhances data reliability for clinical research, potentially improving research reproducibility and facilitating meta-analyses across studies. Healthcare organizations can expect substantial time and resource savings through this one-time comprehensive harmonization process, as harmonized data can be reused seamlessly in different analyses without repeated standardization work. Furthermore, this methodology contributes to the broader goal of healthcare data interoperability, supporting more effective data exchange between systems and institutions while maintaining semantic integrity of clinical measurements.
Future research will focus on three key areas: (1) extending the framework to accommodate diverse data structures across multiple clinical databases, improving its cross-platform applicability; (2) enhancing the system’s contextual understanding through domain-specific improvements to both the embeddings and re-ranking components; and (3) streamlining the validation workflow through improved user interfaces and synthetic training data generation. These advancements will contribute to a more robust, adaptable harmonization system that can meet the evolving needs of clinical data management across healthcare ecosystems.
Acknowledgments
We would like to extend our heartfelt thanks to the open-source community. Their culture of collaboration and shared knowledge is invaluable, contributing not only to this work but to the collective progress of our society. Without the dedication of countless individuals and projects, this work would not have been possible.