### v2

# Scalable Unit Harmonization in Medical Informatics via Bayesian-Optimized Retrieval and Transformer-Based Re-ranking

Abstract
Objective: To develop and evaluate a scalable methodology for harmonizing inconsistent units in large-scale clinical datasets, addressing a key barrier to data interoperability.
Materials and Methods: We designed a novel unit harmonization system combining BM25, sentence embeddings, Bayesian optimization, and a bidirectional transformer based binary classifier for retrieving and matching laboratory test entries. The system was evaluated using the Optum Clinformatics Datamart dataset (7.5 billion entries). We implemented a multi-stage pipeline: filtering, identification, harmonization proposal generation, automated re-ranking, and manual validation. Performance was assessed using Mean Reciprocal Rank (MRR) and other standard information retrieval metrics.
Results: Our hybrid retrieval approach combining BM25 and sentence embeddings (MRR: 0.8833) significantly outperformed both lexical-only (MRR: 0.7985) and embedding-only (MRR: 0.5277) approaches. The transformer-based reranker further improved performance (absolute MRR improvement: 0.10), bringing the final system MRR to 0.9833. The system achieved 83.39% precision at rank 1 and 94.66% recall at rank 5.
Discussion: The hybrid architecture effectively leverages the complementary strengths of lexical and semantic approaches. The reranker addresses cases where initial retrieval components make errors due to complex semantic relationships in medical terminology.
Conclusion: Our framework provides an efficient, scalable solution for unit harmonization in clinical datasets, reducing manual effort while improving accuracy. Once harmonized, data can be reused seamlessly in different analyses, ensuring consistency across healthcare systems and enabling more reliable multi-institutional studies and meta-analyses.
1 Background and Significance
Modern scientific research increasingly depends on the integration of heterogeneous data sources, with inconsistent units of measurement posing a persistent challenge, particularly in laboratory analyses where identical quantities may be reported using diverse conventions. Data harmonization is a critical prerequisite for large-scale clinical research, ensuring interoperability across electronic health record systems and supporting reliable, reproducible medical analyses. Laboratory unit harmonization is especially complex due to domain-specific terminology and the clinical importance of precise unit conversion. Traditional methods for unit harmonization often rely on manually curated rule-based systems and mapping tables developed by domain experts [1]. While these approaches can be effective within narrowly defined contexts, they are not easily scalable and require ongoing maintenance to accommodate emerging laboratory codes and units. This labor-intensive process introduces bottlenecks in research workflows, delays analysis, and increases resource burdens. Emerging methodologies that combine machine learning with information retrieval techniques offer promising solutions to these challenges. Techniques such as sentence embeddings, bidirectional transformers, and optimized retrieval models have demonstrated potential for automating and scaling harmonization processes with improved accuracy and efficiency [2, 3, 4]. In this paper, we present a novel automated unit harmonization system that advances current practices by integrating Bayesian-optimized BM25, sentence embeddings, and transformer-based re-ranking mechanisms. Our system directly addresses major limitations in existing frameworks, including poor scalability to billions of records, limited adaptability to diverse naming conventions, inadequate contextual understanding of laboratory codes, lack of dynamic feedback loops, and overreliance on manual validation. By mitigating these constraints, our approach facilitates seamless data integration, reduces error rates, and accelerates the pace of scientific discovery, ultimately empowering clinical researchers to derive more robust, timely, and reproducible insights from increasingly complex and voluminous healthcare datasets.
2 Materials and Methods
2.1 Datasets
2.1.1 Primary Dataset
The Optum Clinformatics Data Mart (CDM) [5] constitutes our primary evaluation dataset. This comprehensive, de-identified database integrates administrative health claims data from over 84 million individuals across all 50 U.S. states. The resource contains more than 7.5 billion medical and pharmacy claims, documenting healthcare utilization and associated costs. Data elements include member demographics, detailed medical and pharmacy claims, laboratory results, inpatient confinement records, and provider information.
The large scale and heterogeneity of the dataset provide a rigorous testbed for evaluating the effectiveness and scalability of our unit harmonization system. The vast diversity of lab tests, samples, and reported units reflects real-world variability and complexity in clinical data sources, posing significant challenges for accurate and automated harmonization. Furthermore, the inclusion of data from diverse healthcare settings across all U.S. states ensures that the system is evaluated against a wide range of measurement practices and data quality variations.
Given that Optum CDM is extensively used for pharmacoepidemiology, outcomes research, and healthcare analytics, successful harmonization of units within this dataset directly contributes to improving the reliability and consistency of downstream clinical analyses and decision-making. Therefore, demonstrating our system’s performance on this dataset not only validates its technical robustness but also its practical utility in large-scale medical informatics applications.
2.1.2 Reference Dataset
We utilized an internal Labcodes Standard Database for unit harmonization, providing a comprehensive mapping of laboratory tests with standardized information. The database facilitates consistent data representation across diverse sources by linking original units to preferred units and providing necessary conversion factors.
The harmonization process is mapped against multiple fields within this database, including Test Name, Test Label, Synonym, Sample Name, Labcode, Preferred Unit, Actual Unit, Conversion Factor, and various CDISC standardization fields. Although all fields contribute to a comprehensive test definition, harmonization can often be achieved with a minimum of Test Name, Sample Name, and Unit. Inclusion of additional fields improves the accuracy and specificity of the match.
2.2 System Architecture
Our harmonization framework implements a multi-stage pipeline comprising data processing, predictive modeling, and validation components. A key design principle is the modular decomposition of this framework into independent blocks, allowing for focused development, rigorous testing, and iterative improvement of individual components.
2.2.1 Overall Pipeline Structure
The system is organized as an end-to-end processing pipeline (Fig. 1) comprising three main stages: preprocessing, harmonization, and post-processing.
The preprocessing stage begins by retrieving raw data from the source system, Optum Clinformatics Data Mart (2024 Q1, Q2, and Q3). To support downstream harmonization, key fields are extracted, including LOINC codes, units of measurement, frequency of occurrence, and descriptive statistics (minimum, maximum, mean, and standard deviation). This process yields approximately 30,000 unique (test, sample, unit) triads requiring standardization, of which 17,244 were found to have candidate matches in our internal database.
Next, the extracted data undergoes cleaning and normalization to ensure consistency in structure and values. A validation step then confirms the integrity of the LOINC codes and associated metadata. Finally, the dataset is enriched by mapping valid LOINC codes to standardized labels and attributes, providing essential context for the harmonization phase.
Figure 1: Overview of the harmonization pipeline showing the sequential flow of data from raw database entries to validated harmonized units
During harmonization, the system constructs tailored queries that generate ranked proposals for matching equivalents. These queries leverage the hybrid retrieval engine presented in this paper, enabling efficient searches for equivalent data elements. The retrieved results are then processed and structured for evaluation, followed by a reranking step that prioritizes candidates based on relevance scores to enhance overall accuracy.
The pipeline concludes with post-processing, where domain experts review the harmonization outcomes through a custom interface to ensure clinical safety and correctness. Finally, validated results are integrated into the target repository, completing the harmonization process.
As illustrated in Fig. 1, this comprehensive workflow balances automated methods with expert oversight to deliver reliable and high-quality data harmonization.
2.2.2 Predictive Model Architecture
The core intelligence of the system is embodied in the predictive model architecture (Fig. 2), which consists of a Hybrid Retrieval Engine and a Retrieval Reranker.
The Hybrid Retrieval Engine combines two complementary retrieval methods: the Lexical Retrieval Module, which uses the BM25 algorithm for precise keyword matching and excels at exact and fuzzy matches of laboratory terminology; and the Semantic Retrieval Module, which employs sentence embeddings to capture contextual relationships, enabling detection of semantically equivalent terms even when lexical overlap is limited.
A key design feature within this engine is the Bayesian Score Optimizer, which uses Gaussian Process-based optimization to determine the optimal weighting between lexical and semantic scores. This optimization is performed offline before inference, ensuring that precomputed weights guide the weighted combination of retrieval scores during query execution.
Figure 2: Components of the predictive model architecture showing the interaction between retrieval mechanisms and contextual reranking
The Retrieval Reranker serves as the final decision layer. This bidirectional transformer model [6, 7, 8] reevaluates and reorders candidate matches based on deeper contextual understanding, addressing subtle semantic distinctions critical in complex medical terminology.
The retrieval engine relies on an Elasticsearch backend, supporting scalable and flexible querying via JSON schema definitions that govern indexing and search of laboratory metadata. Candidate ranking scores are calculated as a weighted linear combination of BM25 lexical scores and semantic similarity scores measured by cosine distance between embeddings, with weights established by the Bayesian optimization design embedded within the Hybrid Retrieval Engine.
This modular architecture achieves both high accuracy and scalability. Broad candidate coverage is provided by the hybrid retrieval engine, while the reranker ensures precision through deeper semantic analysis. Each component can be independently optimized and updated, maintaining backward compatibility and allowing incorporation of future advances.
2.2.3 Baseline Retrieval Using BM25
Our retrieval system incorporates as a base the BM25 algorithm [9], a state-of-the-art probabilistic relevance framework. BM25 offers a refined modeling of term frequency, along with document length normalization to mitigate frequency bias. The BM25 score for a document $d$ with respect to a query $q$ is computed as:
$\text{score}(d,q)=\sum_{t\in q}\text{IDF}(t)\cdot\frac{\text{TF}(t,d)\cdot(k_{1}+1)}{\text{TF}(t,d)+k_{1}\cdot\left(1-b+b\cdot\frac{|d|}{\text{avgdl}}\right)}$
where:
  • $\text{TF}(t,d)$ is the term frequency of term $t$ in document $d$
  • $\text{IDF}(t)$ is the inverse document frequency of term $t$
  • $|d|$ is the length of document $d$ in words
  • avgdl is the average document length in the collection
  • $k_{1}$ (typically 1.2-2.0) and $b$ (typically 0.75) are free parameters that control term frequency scaling and document length normalization respectively
Our system builds on the strengths of the BM25 ranking function to support both exact and fuzzy matching, optimizing retrieval performance across varying levels of lexical precision. In scenarios requiring exact matching, BM25 prioritizes documents that contain query terms with higher frequency and in close proximity. This capability is particularly beneficial in the context of laboratory code harmonization, where precise terminology—such as LOINC codes or standardized unit expressions—is critical for accurate identification and mapping.
In addition to exact matches, the system supports fuzzy matching to accommodate natural variations in terminology, including differences in spelling, formatting, or word choice. BM25 introduces a graded penalty for such partial matches, allowing relevant documents with similar but non-identical terms to be retrieved while preserving ranking quality. This is especially important in healthcare data, where naming conventions can vary across institutions or over time.
To further enhance retrieval effectiveness, the system incorporates curated synonym lists that expand the query space by mapping equivalent or related terms. This extension is particularly valuable in laboratory contexts, where a single concept may be referred to using multiple terminologies. By integrating synonym expansion into the BM25 framework, the system ensures comprehensive coverage and improves recall without sacrificing precision. Collectively, these enhancements enable the retrieval engine to support robust, flexible, and clinically meaningful harmonization of laboratory units.
2.2.4 Contextual Retrieval via Sentence Embeddings
Sentence embeddings [10, 11, 12] serve as the semantic backbone of our retrieval system, enabling context-aware matching between queries and candidate records that go beyond lexical overlap. These embeddings are dense vector representations of text, where semantically similar sentences are mapped to nearby points in the embedding space. In the context of laboratory terminology, this is crucial for recognizing variant expressions of the same clinical concept, such as “blood urea nitrogen” and “BUN,” or for disambiguating polysemous terms based on surrounding context.
2.2.5 Hybrid Lexical–Semantic Retrieval and Ranking
Our hybrid search system integrates both lexical and semantic search capabilities through a multi-tiered (Fig. 2) designed to enhance retrieval accuracy and contextual relevance. The first component of this architecture involves the generation of combined text embeddings that represent laboratory test descriptions alongside relevant metadata. These embeddings enable the system to capture nuanced semantic relationships between queries and candidate entries, going beyond surface-level term matching.
To refine retrieval precision, the system employs field-specific search boosts that reflect the relative importance of different attributes. For instance, fields such as standardized test names or unit designations may carry more weight in the matching process than auxiliary metadata. These boost factors are configurable and allow the system to dynamically adjust the influence of each field depending on the context of the search.
Structured queries are constructed to leverage both exact and fuzzy matching mechanisms, alongside semantic vector similarity. The system utilizes a custom-built module that parses incoming queries using field-aware syntax, applies the predefined boost factors, and computes a composite score for each candidate result. This score is derived through a weighted combination of BM25 lexical scores and cosine similarity from semantic embeddings, allowing for a balanced integration of symbolic and contextual retrieval. By merging these complementary strategies, the system achieves high recall and precision in laboratory unit harmonization tasks, even in the presence of terminological variation and incomplete data.
2.2.6 Tuning Lexical and Semantic Weights via Bayesian Optimization
To determine the optimal combination of retrieval methods, we implemented Bayesian optimization [13] to fine-tune the weights assigned to various components of the search process. This approach allows for the systematic exploration of the search parameter space, ensuring that the retrieval system operates at peak efficiency while balancing competing factors. The optimization process primarily focused on adjusting field boost weights, which control the relative importance of different fields (e.g., test descriptions or units) in the retrieval process. By tuning these parameters, the system can adapt to the specific needs of each search context.
Another key aspect of the optimization was balancing the contributions from the lexical (BM25) and semantic (embedding-based) search components. Both approaches offer unique advantages: lexical search excels at exact term matching, while semantic search captures deeper contextual relationships. The optimization process sought to find the best weighting between these two components to maximize retrieval effectiveness across diverse queries and data types.
Finally, the optimization aimed to maximize the Mean Reciprocal Rank (MRR) [14], a widely used performance metric that evaluates the ranking quality of retrieved results. By focusing on MRR, the system ensures that relevant harmonization candidates appear near the top of the result list, enhancing the user experience and enabling faster, more accurate decision-making in laboratory unit harmonization tasks. Through this optimization process, we were able to fine-tune the retrieval system to deliver the most relevant results while minimizing errors and inconsistencies.
2.3 Transformer-Based Reranking of Retrieved Candidates
The reranker module is responsible for ranking candidate harmonizations presented by previous modules by evaluating their semantic compatibility with the input laboratory test. It operates on test triads, each comprising a test name, a sample type, and a measurement unit. For each candidate triad, the reranker computes a compatibility score based on its similarity to the input, enabling the system to prioritize the most plausible harmonization suggestions. This component plays a critical role in filtering out false matches and refining the output quality of the harmonization pipeline.
To address the complexity of laboratory data, which often includes specialized terminology, institution-specific conventions, and heterogeneous representations, the reranker incorporates tokenization and handling mechanisms specifically adapted to biomedical language. This enables the model to more effectively understand and compare test descriptions that differ in surface form but refer to the same concept.
The reranker is further enhanced by contrastive learning objectives that help the model distinguish between closely related yet non-equivalent test triads. This is especially important for laboratory terms that may differ only slightly in wording or structure but convey distinct clinical meanings. By leveraging a pairwise comparison strategy, the reranker can evaluate subtle differences and improve ranking precision.
Importantly, the model is designed to support incremental learning. As the system is used in real-world settings, manually verified harmonizations are incorporated back into the training set. This continuous learning process allows the reranker to evolve over time and adapt to emerging terminology and use cases, improving its long-term performance and generalization capabilities.
2.4 System Performance Evaluation
We evaluated the performance of our approach through a comprehensive suite of metrics, ensuring that both the retrieval and re-ranking components of the system were thoroughly assessed. For retrieval performance, we measured metrics such as Precision@k, Recall@k, and Mean Reciprocal Rank (MRR). These metrics allow us to evaluate how well the system retrieves relevant candidates across different ranks, ensuring that the most relevant harmonization suggestions appear early in the result set and assessing the system’s overall retrieval effectiveness.
The re-ranking performance was evaluated using standard classification metrics, including Accuracy, Precision, Recall, and the F1 score. These metrics provide insights into the effectiveness of the reranker in classifying and ranking candidate harmonizations. By analyzing these performance indicators, we could assess how well the reranker distinguishes between relevant and irrelevant harmonization proposals, and how balanced its classification performance is across different types of test descriptions.
To assess the scalability of the system, we measured key factors such as indexing throughput, query latency, batch processing efficiency, and memory usage. These metrics are critical for ensuring that the system can handle large datasets and operate efficiently at scale, especially when processing millions of laboratory records.
Additionally, the manual validation process was tracked through a custom tagging system, which categorized each harmonization decision into one of the following statuses: ”Missing,” ”Verified,” ”Pending,” ”Human,” ”Copy,” or ”Reranked”. This system enabled us to monitor the progress and quality of harmonization proposals, track user interactions with the system, and ensure that manual corrections were incorporated into the system’s learning process, contributing to its ongoing refinement.
3 Results
In this section, we present the results from the different modules that comprise our retrieval system. We begin with the selection of core components, including BM25 parameter settings and sentence embedding models. This is followed by Bayesian optimization of the weighting parameters used in the combined re-ranking module, training of the re-ranker for optimal ranking performance, and finally, a comparison of the overall system performance under the optimized configuration versus simpler baselines.
3.1 BM25 Parameter Selection
We use the default BM25 configuration provided by Elasticsearch for our retrieval experiments. BM25 is a widely adopted probabilistic ranking function that scores documents based on term frequency, inverse document frequency, and document length normalization. Elasticsearch’s implementation includes two key parameters: $k_{1}$, which controls term frequency saturation and is set to 1.2 by default, and $b$, which governs the degree of document length normalization, with a default value of 0.75. These parameters balance the influence of term frequency and document length on the relevance score. We retained the default settings, as they are generally effective across a range of domains and retrieval tasks, and preliminary experiments did not indicate significant performance gains from additional tuning.
3.2 Sentence Embedding Selection
To select an appropriate sentence encoder for our retrieval task involving laboratory tests, samples, and unit combinations, we evaluated several pre-trained models based on their performance in the Massive Text Embedding Benchmark (MTEB), as shown in Table 1. Our primary selection criteria included model size—favoring models with fewer than one billion parameters for efficiency—and the overall mean task score, which reflects performance across a diverse set of semantic tasks. This balance ensured both strong generalization capabilities and practical resource requirements for integration into our system.
Model # Parameters Embedding Dim. Max Tokens Mean (Task)
SFR-Embedding-Mistral 7B 4096 32768 60.90
stella_en_1.5B_v5 1.5B 8960 131072 56.53
NV-Embed-v2 7B 4096 32768 56.29
stella_en_400M_v5 435M 4096 8192 48.32
bge-small-en-v1.5 33M 512 512 43.76
all-MiniLM-L6-v2 22M 384 256 41.39
Table 1: Comparison of embedding models on MTEB Mean (Task) score.
Among the evaluated models meeting the size restriction, stella_en_400M_v5 — a transformer with approximately 435 million parameters trained using Matryoshka Representation Learning (MRL) [15] — demonstrated strong general performance. While this model produces 4096-dimensional embeddings, the Matryoshka training method distributes semantic information hierarchically across the embedding dimensions, allowing effective selection of lower-dimensional subsets without significant loss of information. Leveraging this property, we used 1024-dimensional representations for our experiments, balancing embedding compactness with semantic richness.
In comparison, lighter models such as bge-small-en-v1.5 and all-MiniLM-L6-v2, with 33 million and 22 million parameters respectively, showed lower Mean (Task) scores, reflecting a trade-off between model size and embedding quality.
Our approach uses these embedding models to generate a candidate set of matches based on semantic similarity. The responsibility of fine-grained selection according to the particularities of our application is delegated to a reranker module, which applies domain-specific criteria to reorder the candidates for the final retrieval.
We also explored fine-tuning using a synthetic dataset of sentence pairs labeled as semantically equivalent or dissimilar, focusing on laboratory-specific terminology. However, our experiments showed that general-purpose pre-trained models consistently outperformed these fine-tuned variants, indicating that large-scale general-domain training still provides strong out-of-the-box performance. Future work may revisit fine-tuning using alternative strategies or additional domain supervision.
3.3 Bayesian Optimization of BM25 and Sentence Embedding Parameters
Figure 3: Bayesian optimization convergence showing the progression of MRR optimization over iterations. The plot displays the objective function values (MRR) across optimization iterations.
The Bayesian optimization process, shown in Figure 3, was performed over approximately 120 different parameter configurations using a validation set of 2,500 distinct laboratory codes. The optimization balanced the lexical contributions of the TEST, SAMPLE, and UNIT fields and a semantic component based on cosine similarity. The cosine similarity was clipped at zero using a max(0, cosine similarity) function before being weighted—assigning zero when dissimilar and up to the full optimized weight when similar. The objective function optimized was the Mean Reciprocal Rank (MRR) across the validation set.
3.4 Reranker Implementation and Training
Figure 4: Validation curves of the training process for the transformer-based reranker. The plot shows the progression of validation loss and validation accuracy over training epochs.
The reranker is implemented using a fine-tuned version of RoBERTa-base-PM-M3-Voc-train-longer model [16], a biomedical variant of RoBERTa [8] that was pre-trained on PubMed abstracts, PMC full-text articles, and MIMIC-III clinical notes. This model employs a byte-pair encoding vocabulary derived from PubMed content and was further pre-trained for 50,000 additional steps to improve domain-specific understanding. Its adaptation to clinical and biomedical language makes it particularly suitable for evaluating laboratory test descriptions.
To fine-tune the model for the harmonization task, we constructed a synthetic dataset composed of 3,135,557 labeled pairs of laboratory test triads. Each pair contains two triads, and the label indicates whether the two triads are semantically compatible or not. The dataset includes pairs with varying levels of dissimilarity to simulate real-world ambiguity. Some pairs differ in all three components—test, sample, and unit—while others differ in only one or two. Pairs labeled as compatible share the same meaning across all components but use different lexical expressions, such as synonyms or variant naming conventions. This stratified construction ensures the model learns to recognize not only exact matches but also semantically equivalent formulations. The reranker finetuning process using as training a synthetic dataset is illustrated in Figure 4.
Training is conducted using binary cross-entropy loss, with the model learning to classify whether a given pair of triads represents a valid equivalence. The output probability serves as a compatibility score and is used directly as the ranking metric. Additionally, contrastive learning objectives are incorporated during training to sharpen the model’s ability to differentiate between closely related but non-equivalent pairs. This helps improve the reranker’s performance in borderline cases where semantic distinctions are subtle.
3.5 Performance evaluation
We evaluated our harmonization approach on a labeled dataset of 17,243 queries from the Optum database, employing the optimal parameters identified through the prior Bayesian optimization. In the following sections, we present the results of this performance evaluation.
3.5.1 Comparative Approach Analysis
We conducted three controlled experiments to evaluate the performance of distinct retrieval paradigms, each aimed at highlighting the strengths of specific retrieval techniques.
The first experiment, Lexical-Only Retrieval, implemented a fully optimized lexical retrieval pipeline that combined BM25, field-specific weighting, fuzzy string matching, and synonym expansion. This approach focused solely on exact term matching and lexical features. The results from this experiment showed a best MRR of 0.7985, highlighting the effectiveness of lexical retrieval in terms of precision and matching the exact terms present in the query.
The second experiment, Embedding-Only Retrieval, evaluated the retrieval performance using exclusively sentence-level semantic similarity, derived from a general-purpose sentence encoder. This approach relied entirely on semantic embeddings to capture deeper contextual relationships between queries and candidate results. The best MRR in this case was 0.5277, demonstrating the ability of semantic embeddings to capture broader relationships but also revealing their limitations in terms of precise domain-specific matching.
The third experiment, Hybrid Retrieval, combined both semantic and lexical signals within a unified retrieval framework. This approach synthesized the strengths of both methods, using lexical signals for precision and embeddings for semantic coverage. The best MRR achieved by this hybrid approach was 0.8833, showcasing a significant improvement over the individual approaches. These results underscore the complementary nature of semantic and lexical retrieval, where embeddings enhance coverage and recall by capturing semantic relationships beyond exact keyword matching, while lexical signals provide precision and discriminative power for domain-specific terminology.
Together, these experiments demonstrate that a hybrid retrieval approach, which combines both semantic and lexical features, leads to the most effective retrieval performance, offering a balanced solution that maximizes both recall and precision.
3.5.2 Overall Performance Metrics
Table 2 presents comprehensive evaluation metrics for our hybrid retrieval model:
Table 2: Comprehensive evaluation metrics for the hybrid retrieval model
Metric Hybrid Model Value
Mean Reciprocal Rank (MRR) 0.8833
Mean Average Precision (MAP) 0.8131
Precision@10 0.5863
Recall@10 0.9700
NDCG@10 0.5291
Success@10 0.9700
MRR@10 0.8833
Queries Evaluated 17,243
Queries With Results 17,243 (100%)
The identical values of Recall@10 and Success@10 (both 0.9700) stem from our evaluation setup, which assumes exactly one relevant result per query.
3.5.3 Performance at Different Cutoff Thresholds
Table 3 shows retrieval metrics at varying rank cutoffs:
Table 3: Retrieval metrics at varying rank cutoffs
Metric k=1 k=3 k=5
Precision@k 0.8339 0.7116 0.6520
Recall@k 0.8339 0.9224 0.9466
NDCG@k 0.8339 0.5992 0.5353
Success@k 0.8339 0.9224 0.9466
MRR@k 0.8339 0.8746 0.8801
These metrics highlight the behavior of the retrieval system as we increase the cutoff threshold $k$:
- Recall@k and Success@k increase with larger $k$ Precision@k and NDCG@k decrease with increasing $k$ MRR@k increases slightly with larger $k$, eventually saturating
3.5.4 Reranker Performance
The transformer-based reranker was implemented as a post-processing stage after initial hybrid retrieval. The reranker selectively overrides the initial ranking when its confidence score exceeds that of the top-1 entry from the hybrid retrieval phase.
Absolute MRR Improvement: 0.10
This improvement represents a substantial enhancement over the already strong hybrid retrieval model, bringing the final system MRR to 0.9833. The reranker was particularly effective at correcting cases where lexically similar but semantically different terms were initially ranked higher than the correct match.
4 Discussion
The experimental results provide strong evidence for the effectiveness of our hybrid harmonization approach in medical terminology retrieval. The substantial performance gap between the hybrid model (MRR: 0.8833) and both the lexical-only (MRR: 0.7985) and embedding-only (MRR: 0.5277) variants confirms that these retrieval paradigms capture complementary aspects of query-document relevance.
The relatively poor performance of the embedding-only approach suggests that while general-purpose sentence encoders capture semantic relatedness, they may lack the specificity required for precise medical terminology matching. Conversely, the lexical approach demonstrates strong discriminative power but may miss semantically equivalent expressions with limited lexical overlap.
Our hybrid architecture effectively addresses these limitations by leveraging the complementary strengths of both approaches. The performance metrics across different rank cutoffs demonstrate that the system achieves an optimal balance between precision and recall, with over $83\%$ of queries returning the correct result in the top position.
The addition of the transformer-based reranker as a final stage provides a critical refinement layer, addressing cases where the initial retrieval components make errors due to complex semantic relationships or domain-specific nuances in medical terminology. The significant improvement in MRR achieved by the reranker underscores the value of deep contextual understanding in medical term harmonization tasks.
Existing harmonization systems encounter several critical challenges that hinder their effectiveness in real-world clinical settings. One major limitation is the limited scalability of these systems, making it difficult to process the vast volumes of laboratory records required for large-scale clinical research. As the volume of data continues to grow, traditional methods struggle to handle the increasing demand for faster and more efficient harmonization.
Another significant challenge is the insufficient handling of variations in naming conventions and abbreviations. Laboratory codes and terminologies often vary across different institutions and datasets, and existing systems are not always equipped to manage these discrepancies effectively. This leads to inconsistencies in data representation, making it harder to perform accurate and reliable harmonization.
Furthermore, many harmonization systems fail to incorporate contextual information when matching laboratory codes. Without context, these systems may misinterpret or incorrectly align terms that appear similar but have different meanings depending on the context, which compromises the quality and accuracy of the harmonization process.
A related issue is the lack of efficient feedback mechanisms for continuous performance improvement. As laboratory terminologies evolve and new codes emerge, existing systems often do not have an effective way to incorporate feedback or adapt to changes in real-time, leading to stagnation and a failure to keep pace with evolving practices.
Finally, there is a high dependency on manual curation and validation in many current harmonization approaches. Manual intervention is time-consuming, prone to human error, and often not scalable, making it difficult to maintain quality and consistency as datasets grow. These challenges highlight the need for more automated, scalable, and adaptive harmonization systems in clinical settings.
Our system addresses these limitations through its modular architecture, hybrid retrieval approach, and reranker component. The Bayesian optimization of weightings between lexical and semantic components allows the system to adapt to the specific characteristics of medical terminology, while the reranker provides deeper contextual understanding.
The system’s performance on the large-scale Optum dataset (7.5 billion entries) demonstrates its scalability and effectiveness in real-world settings. By automating much of the harmonization process, it reduces the manual effort required while maintaining high accuracy.
5 Conclusion
We presented a scalable and efficient framework for unit harmonization in clinical datasets using a combination of BM25, sentence embeddings, a reranker based on bidirectional transformers, and Bayesian optimization techniques. Our system automates the harmonization process, reducing manual effort and improving accuracy. The results obtained from the Optum Clinformatics Data Mart dataset demonstrate that the methodology is effective and adaptable to large datasets, making it a promising solution for future healthcare data standardization efforts.
The implications of this work extend beyond technical achievements to address fundamental challenges in healthcare data management. By providing a consistent and standardized approach to unit harmonization, our framework significantly enhances data reliability for clinical research, potentially improving research reproducibility and facilitating meta-analyses across studies. Healthcare organizations can expect substantial time and resource savings through this one-time comprehensive harmonization process, as harmonized data can be reused seamlessly in different analyses without repeated standardization work. Furthermore, this methodology contributes to the broader goal of healthcare data interoperability, supporting more effective data exchange between systems and institutions while maintaining semantic integrity of clinical measurements.
Future research will focus on three key areas: (1) extending the framework to accommodate diverse data structures across multiple clinical databases, improving its cross-platform applicability; (2) enhancing the system’s contextual understanding through domain-specific improvements to both the embeddings and re-ranking components; and (3) streamlining the validation workflow through improved user interfaces and synthetic training data generation. These advancements will contribute to a more robust, adaptable harmonization system that can meet the evolving needs of clinical data management across healthcare ecosystems.
Acknowledgments*
We would like to extend our heartfelt thanks to the open-source community. Their culture of collaboration and shared knowledge is invaluable, contributing not only to this work but to the collective progress of our society. Without the dedication of countless individuals and projects, this work would not have been possible.